OpenStudy (anonymous):

Please help f(x) = -8 cos ( sin ( x ^{4} )) find f'(x)

OpenStudy (anonymous):

Okay so im just going to identify composite functions: \[f(x)=-8\cos[\sin(x^4)]=-8\cos[\sin(g(x))]=-8\cos[h(x)]=-8j(x)\] So then we have simplified the question to: \[\eqalign{ \frac{d}{dx}f(x)&=[-8j(x)]' \\ &=-8j'(x) \\ &=-8[\cos(h(x))]' \\ &=-8[-\sin(h(x))h'(x)] \\ &=8\sin[h(x)]h'(x) \\ &=8\sin[\sin(g(x))][\sin(g(x))]' \\ &=8\sin[\sin(x^4)][\cos(g(x))g'(x)] \\ &=8\sin[\sin(x^4)]\cos[g(x)]g'(x) \\ &=8\sin[\sin(x^4)]\cos(x^4)(x^4)' \\ &=8\sin[\sin(x^4)]\cos(x^4)(4x^3) \\ &=32x^3\sin[\sin(x^4)]\cos(x^4) \\ }\]

OpenStudy (anonymous):

Thanks, that was helpful

OpenStudy (anonymous):

No prob!